OpenStudy (caozeyuan):

a curve is defined as x=cos(t)+t*sin(t),y=sin(t)-t*cos(t), FInd the area of surface of revolution if t is between 0 and a half pi, and the curve is rotated around the x-axis

OpenStudy (amistre64):

define (x')^2 and (y')^2

OpenStudy (caozeyuan):

dx / dt = t cos t, dy / dt = t sin t That's what the mark scheme said

OpenStudy (amistre64):

^2 them and take the sqrt

OpenStudy (amistre64):

should end up with t i think

OpenStudy (caozeyuan):

but the derivatives are not right, are they?

OpenStudy (amistre64):

x = cos(t)+t sin(t) x' = -sin + sin + t cos y = sin(t)-t cos(t) y' = cos - cos + t sin i dont see why they wouldnt be

OpenStudy (amistre64):

t^2(cos^2+sin^2) = t^2 and sqrt(t^2) = t \[2\pi\int yt~ dt\]

OpenStudy (caozeyuan):

OMG!!!! My mistake ! I just get the derivative wrong! I'm soooooo stupid!

OpenStudy (amistre64):

practice helps, and then we still make mistakes at times :)

OpenStudy (caozeyuan):

Now I need to evaluate \[\int\limits_{?}^{?} t ^{2} \cos t\]

OpenStudy (caozeyuan):

How to do it ? Confused

OpenStudy (amistre64):

by parts du int v ------------ cos t + t^2 sin t - 2t -cos t + 2 -sin t - 0 so... t^2 sin t + 2t cos t - 2 sin t